### v1

@@ 2 First-Order Electroweak Phase Transition in the 2HDM
This relates the vacuum expectation value of the Higgs field at the percolation stage, $v_{p}$, to the percolation temperature $T_{p}$. The parameter region characterised by such quantities can be considered sufficiently protected against baryon number washout at a stage in the thermal history of the universe when around a third of the comoving volume has been converted to the broken electroweak phase. [^1] For an overview of related literature, cf. e.g. the recent review on the relation between particle physics, SFOEWPTs and gravitational waves Athron:2023xlk .
@@
Currently, all signal strength measurements of the discovered Higgs boson are consistent with the SM prediction within the experimental and theoretical uncertainties ATLAS:2022vkf ; CMS:2022dwd . Therefore, any viable BSM model with an extended Higgs sector must be able to accommodate an SM-like Higgs boson. Within the 2HDM, assuming that the SM-like Higgs boson is identified with the state $h$, there are two possibilities to achieve this Gunion:2002zf ; Carena:2013ooa ; Bernon:2015qea . The first one is by considering the limit $\cos\!\left(\beta-\alpha\right)\to 0$ (or analogously $\beta-\alpha\to\pi/2$), since, in this limit, the tree-level couplings of $h$ to the SM particles yield the SM prediction, i.e. $\zeta_{V}^{h}=\zeta_{f}^{h}=1$ (see Eqs. (2) and (4)). [^3] In this limit, the $h$ triple and quartic self-couplings also have the SM value at tree level. This is often referred to as alignment without decoupling, or simply alignment limit, since the masses of the other BSM Higgs bosons can, in general, be light. The second option is to make the masses of the unobserved Higgs bosons very heavy, namely $m_{H},\,m_{A},\,m_{H^{\pm}}\gg v$, while keeping the quartic couplings in the potential at ($\order{1}$ ). With these considerations, the alignment limit is obtained ($\cos\!\left(\beta-\alpha\right)\to 0$) together with the condition that $\bar{m}\sim m_{H},\,m_{A},\,m_{H^{\pm}}\gg v$. This ensures that the scalar interactions between Higgs bosons do not increase with their masses, keeping their contributions to physical observables negligible. This is known as alignment with decoupling, or alternatively decoupling limit. While in the alignment limit, BSM bosons could, in principle, induce sizable radiative corrections (or, in other words, give rise to non-decoupling effects), in the decoupling limit, these corrections vanish, and the SM predictions are always recovered. This distinction between alignment and decoupling is crucial because the existence of an SFOEWPT in the 2HDM requires large radiative corrections in order to induce a potential barrier, and therefore it is incompatible with the decoupling limit (see, for instance, Refs. Dorsch:2014qja ; Dorsch:2016nrg ; Basler:2016obg ; Dorsch:2017nza ; Biekotter:2023eil and references therein).
@@
We employ the public code BSMPTv3 Basler:2024aaf to compute the above potential, which includes, at zero temperature, the tree-level potential $V_{\mathrm{tree}}$ of the 2HDM, the one-loop Coleman-Weinberg effective potential $V_{\mathrm{CW}}$ Coleman:1973jx , and a finite counterterm potential $V_{\mathrm{CT}}$, such that an on-shell-like scheme is adopted to fix the mixing matrix elements in the Higgs mass matrices to their tree-level values at zero temperature Basler:2018cwe . In consequence, the one-loop corrected input parameters in Eq. (5) are fixed to their tree-level values at $T=0$. In addition, the thermal corrections to the effective potential $V_{\mathrm{T}}$ are included in the high-temperature limit Quiros:1994dr ; Quiros:1999jp . The temperature-corrected daisy resummation $V_{\mathrm{daisy}}$ is also included following the Arnold-Espinosa method Arnold:1992rz . For more details, we refer the reader to Basler:2018cwe ; Basler:2024aaf .
To explore the possibility of an SFOEWPT within the 2HDM, we have performed a parameter scan of the model, such that the input parameters of the model are varied randomly in the following intervals
@@
while $m_{h}=125.25$ GeV as it is identified with the SM-like Higgs boson. We confront the scan points with the most relevant theoretical and experimental constraints to date as implemented by the public code ScannerS Coimbra:2013qq ; Muhlleitner:2020wwk . More concretely, on the theoretical side, we consider a point to be allowed if it satisfies the requirement of tree-level perturbative unitarity Akeroyd:2000wc ; Ginzburg:2005dt , the potential is bounded from below (at $T=0$Deshpande:1977rw , and if the EW minimum of the potential is the absolute minimum of the potential Barroso:2013awa . From the experimental side, all allowed points are required to lie in the $2\sigma$ uncertainty band of the current determination of the oblique parameters $S,\,T,\,U$ Peskin:1990zt ; Grinstein:1991cd ; Altarelli:1991fk ; Peskin:1991sw ; Burgess:1993vc from the EW fit Haller:2018nnx at the one-loop level (using the formulae from Grimus:2008nb ). Furthermore, we check that all points are allowed by the current bounds from LHC searches for BSM scalars and that the predicted SM-like Higgs boson signal strenghts are within $2\sigma$ with respect to the SM prediction (that is $\chi^{2}-\chi^{2}_{\mathrm{SM}}<6.18$), where we apply these constraints with the help of the public code HiggsTools Bahl:2022igd (formerly known as HiggsBounds Bechtle:2013wla ; Bechtle:2011sb ; Bechtle:2008jh and HiggsSignals Bechtle:2020uwn ; Bechtle:2013xfa ). Finally, we check that the points are allowed at the 95% CL level by present measurements of flavour-changing processes mediated by neutral currents Haller:2018nnx . In order to increase the amount of points allowed after an electroweak precision program at the $Z$ pole (as we will discuss in Sec. 3.1), we performed an additional dedicated scan with $m_{A}\sim m_{H^{\pm}}$ and/or $m_{H}\sim m_{H^{\pm}}$ to ensure that the contributions to the $T$ parameter (or alternatively the $\rho$ parameter) remain small Pomarol:1993mu ; Haber:2010bw .
@@ 3.1 SFOEWPT versus a Precision $Z$ Pole Programme
Figure 1: Projections of the SFOEWPT scan as detailed in the main body of the text. (a) shows the compatibility with the $S$ and $T$ parameters at 68% CL (dotted blue line) and 95% CL (full blue line). Here, light grey points are parameter points compatible with the constraints specified in Sec. 2, dark grey points additionally fulfil the SFOEWPT constraint $\xi_{p}>1$, and colored points fulfil the projected constraints on the EWPOs representatively obtained at a future FCC-ee for its $\sqrt{s}=m_{Z}$ phase. The $Z$-pole programme will efficiently single out a relatively small parameter space of the SFOEWPT-preferred parameter region. (c) highlights the relation of this region in the context of the expected 125 GeV signal strength constraints obtainable at the HL-LHC at 68% CL (dotted red line) and 95% CL (dashed red line), displayed in the $\tan\beta$ versus $\cos(\beta-\alpha)$ plane. (b) and (d) show the contours of the allowed mass spectra after successively applying the constraints.
The results for our parameter scan are presented in Fig. 1. As expected, parameter choices that are relatively far from the alignment limit can be efficiently constrained with a precision $Z$ pole programme at an $e^{+}e^{-}$ machine. Sensitivity here extends well beyond the sensitivity of a coupling analysis for the 125 GeV Higgs boson at the HL-LHC, which is highlighted in Fig. 1(c). [^4] To obtain the confidence intervals of the HL-LHC we performed a $\chi^{2}$ fit to test the projected values of the coupling modifiers to $b$- and $t$-quarks, $W$ and $Z$ bosons, gluons, and $\tau$ and $\mu$ leptons from Ref. CMS:2025hfp againts the 2HDM predictions from Eqs. (2) and (4). The requirement of a significant mass splitting between the heavy neutral scalar and pseudoscalar mass, $m_{H}$ and $m_{A}$, respectively, to achieve an SFOEWPT in the 2HDM (see also Dorsch:2014qja ; Dorsch:2016nrg ; Basler:2016obg ; Dorsch:2017nza ; Biekotter:2023eil ) stands in direct tension with an SM-like outcome of the oblique electroweak precision observables (EWPO) analysis for the expected FCC-ee precision, cf. Fig. 1(b). Points that still admit an SFOEWPT according to Eq. (1) are forced towards the alignment limit $\cos(\beta-\alpha)\simeq 0$, yet at relatively small mass scales, to modify the electroweak cross-over to an SFOEWPT, cf. Fig. 1(d). As already mentioned, signal strength analyses at the HL-LHC do not provide significant constraints for this parameter range. However, the mass scale of the exotic Higgs bosons falls into a range that is well accessible by ATLAS and CMS. We will turn to this in the next section.
@@ 3.2 Context with LHC Higgs Partner Searches
If the exotic states have masses above the $t\bar{t}$ threshold, the heavy scalars $H$ and $A$ predominantly decay into these final states. This holds for the largest part of the successful parameter points of our scan, detailed in Sec. 2. Both ATLAS and CMS are actively searching for heavy scalars in the $t\bar{t}$ channels CMS:2019pzc ; ATLAS:2024vxm . It is well-established that interference effects, both between different Higgs resonances (signal–signal interference) Basler:2019nas ; Bahl:2025you and between the Higgs signal and the QCD background, can significantly distort the expected Breit–Wigner resonance shapes Gaemers:1984sj . In some cases, these effects can even result in a complete cancellation of the signal Jung:2015gta , or severely reduce experimental sensitivity, particularly when the net effect is a depletion rather than an excess of events.
@@
Figure 2: Interference-corrected (signal-signal as well as signal-background) cross sections compared to the pure signal process for LHC production of the uncharged heavy scalars, $H$ (blue points) and $A$ (red points), respectively, as a function of their respective mass, for the parameter points of the scan detailed in Sec. 2.
@@ 3.3.1 Elements of the Calculation
Figure 4: Production of the SM-like Higgs boson in association with a $Z$ boson at the FCC-ee, in relation to the decoupling limit at $\sqrt{s}=240~{}\text{GeV}$ for the scan detailed in Sec. 2. Blue points denote leading-order predictions, which are all compatible with the expected FCC-ee sensitivity of $hZ$ production of $\Delta\sigma/\sigma\simeq 0.4\%$; all points are compatible with an SM outcome of the FCC-ee $Z$-pole programme. Including radiative corrections (red points) modifies these predictions, creating the discovery potential at a lepton collider even when $Z$ pole observables are insensitive. The green region is the experimental sensitivity $\Delta\sigma/\sigma\simeq 0.4\%$. The decoupling cross section $\sigma_{\text{decoup}}$ is evaluated consistently at LO (blue points) and NLO (red points).
@@ 3.3.2 Higgs Production in $e^{+}e^{-}$ Collisions - Discrimination Beyond EWPOs
We are now ready to turn to the implications of our previous results for a precision investigation of associated Higgs production at an $e^{+}e^{-}$ collider. A Higgs machine will typically run at a centre-of-mass energy of $\sqrt{s}=240~{}\text{GeV}$ where the $e^{+}e^{-}\to hZ$ cross section has a maximum. The number of observables that are accessible in this instance is limited compared to hadron machines that probe a plethora of processes across a large range of energy scales. However, when viewed in the context of concrete model predictions, the overall sensitivity that can be expected at an FCC-ee of
$\frac{\Delta\sigma(e^{+}e^{-}\to hZ)}{\sigma(e^{+}e^{-}\to hZ)}\simeq 0.4\%\,,$ (11)
@@
predominantly achieved from an analysis of the $Z$ boson recoil spectrum FCC:2018evy ; Azzurri:2021nmy , carries a very large potential for new physics discovery.
@@
The parameter region of our 2HDM scan that is consistent with the $Z$ pole programme is forced into the 2HDM alignment limit. Hence, it is not surprising that most of the parameter points are consistent with the uncertainty of Eq. (11) at the leading order. Consistency with $S,\,T$ forces the interactions of the SM-like Higgs boson into a region where the leading-order cross section is largely consistent with the SM (the 2HDM cross section in the decoupling limit is around $\sigma_{\mathrm{decoup}}=0.222$ pb at LO and $0.211$ pb at NLO), cf. Fig. 4.
@@
The requirement of an SFOEWPT, however, is incompatible with this decoupling limit (the latter approaches the SM for which the phase transition is a cross-over). Hence, we can expect quantum corrections to be relevant for the points that single out an SFOEWPT, also when they are aligned with the SM expectation of $Z$ pole measurements. Indeed, as shown in Fig. 4, the weak radiative corrections as detailed in Sec. 3.3.1, modify the 2HDM cross sections away from the decoupling limit (the SM). More importantly, the modifications are typically large enough for the $e^{+}e^{-}\to hZ$ programme to detect new physics based on the expectation of Eq. (11).
@@
It is worthwhile mentioning that the discrimination on the basis of the $hZ$ cross section is not in a one-to-one correspondence with the requirement of an SFOEWPT. It is the predictive power of renormalisability that drives experimentally detectable model correlations beyond leading order in a theoretically controlled way. An SFOEWPT is driven by the requirement of a relatively light exotic particle spectrum and a departure from SM-expected Higgs interactions. In any UV-complete scenario (and in particular for the 2HDM considered here), these requirements select a region of the parameter space that is, to a large extent, experimentally distinguishable from the SM in $hZ$ production at a 240 GeV lepton collider. The radiative corrections are not dominantly sensitive to the mass splittings themselves, but to their correlated effects in terms of inter-Higgs couplings and radiative corrections to mixing angles. These effects are two-loop suppressed for the $Z$ pole programme; $hZ$ production accessing these correlation modifications at one-loop level makes them large enough for detection within the expected experimental sensitivity.

### v3

@@ 2 First-Order Electroweak Phase Transition in the 2HDM
This relates the vacuum expectation value of the Higgs field at the percolation stage, $v_{p}$, to the percolation temperature $T_{p}$. The parameter region characterised by such quantities can be considered sufficiently protected against baryon number washout at a stage in the thermal history of the universe when around a third of the comoving volume has been converted to the broken electroweak phase. [^1] For an overview of related literature, cf. e.g. the recent review on the relation between particle physics, SFOEWPTs and gravitational waves Athron:2023xlk . We note that perturbative studies for scenarios with values of $\xi_{p}$ that are closer to the regime where the phase transition is not first order any more, can be misleading in studies of SFOEWPTs from extended scalar sectors, see, e.g. Niemi:2024axp . Our results should therefore be understood with these caveats in mind.
@@
[…]
@@
We employ the public code BSMPTv3 Basler:2024aaf to compute the above potential, which includes, at zero temperature, the tree-level potential $V_{\mathrm{tree}}$ of the 2HDM, the one-loop Coleman-Weinberg effective potential $V_{\mathrm{CW}}$ Coleman:1973jx , and a finite counterterm potential $V_{\mathrm{CT}}$, such that an on-shell-like scheme is adopted to fix the mixing matrix elements in the Higgs mass matrices to their tree-level values at zero temperature Basler:2018cwe . In consequence, the one-loop corrected input parameters in Eq. (5) are fixed to their tree-level values at $T=0$. In addition, the thermal corrections to the effective potential $V_{\mathrm{T}}$ are included in the high-temperature limit Quiros:1994dr ; Quiros:1999jp . The temperature-corrected daisy resummation $V_{\mathrm{daisy}}$ is also included following the Arnold-Espinosa method Arnold:1992rz . For more details, we refer the reader to Refs. Basler:2018cwe ; Basler:2024aaf .
To explore the possibility of an SFOEWPT within the 2HDM, we have performed a parameter scan of the model, such that the input parameters are varied randomly in the following intervals
@@
while $m_{h}=125.25$ GeV as it is identified with the SM-like Higgs boson. We confront the scan points with the most relevant theoretical and experimental constraints to date as implemented by the public code ScannerS Coimbra:2013qq ; Muhlleitner:2020wwk . More concretely, on the theoretical side, we consider a point to be allowed if it satisfies the requirement of tree-level perturbative unitarity Akeroyd:2000wc ; Ginzburg:2005dt , the potential is bounded from below (at T=0T=0Deshpande:1977rw , and if the EW minimum of the potential is the absolute minimum of the potential Barroso:2013awa . From the experimental side, all allowed points are required to lie in the $2\sigma$ uncertainty band of the current determination of the oblique parameters S,T,US,\,T,\,U Peskin:1990zt ; Grinstein:1991cd ; Altarelli:1991fk ; Peskin:1991sw ; Burgess:1993vc from the EW fit Haller:2018nnx at the one-loop level (using the formulae from Grimus:2008nb ). Furthermore, we check that all points are allowed by the current bounds from LHC searches for BSM scalars and that the predicted SM-like Higgs boson signal strengths are within $2\sigma$ with respect to the SM prediction (that is $\chi^{2}-\chi^{2}_{\mathrm{SM}}<6.18$), [^4] This choice is widely used in the literature due to its convenience, since it does not require knowledge of the best fit in a generic model and it facilitates the comparison with other models. This assumes that the SM is close to the best fit of the model (which is reasonable since we have not measured sizable deviations from the SM predictions) and that only the angles $\alpha$ and $\beta$ play an important role in the signal strength predictions. where we apply these constraints with the help of the public code HiggsTools Bahl:2022igd (formerly known as HiggsBounds Bechtle:2013wla ; Bechtle:2011sb ; Bechtle:2008jh and HiggsSignals Bechtle:2020uwn ; Bechtle:2013xfa ). Finally, we check that the points are allowed at the 95% CL level by present measurements of flavour-changing processes mediated by neutral currents Haller:2018nnx . In order to increase the amount of points allowed after an electroweak precision program at the $Z$ pole (as we will discuss in Sec. 3.1), we performed an additional dedicated scan with $m_{A}\sim m_{H^{\pm}}$ and/or $m_{H}\sim m_{H^{\pm}}$ to ensure that the contributions to the $T$ parameter (or alternatively the $\rho$ parameter) remain small Pomarol:1993mu ; Haber:2010bw .
@@ 3.1 SFOEWPT versus a Precision $Z$ Pole Programme
Figure 1: Projections of the SFOEWPT scan as detailed in the main body of the text. Panel 1(a) shows the compatibility with the $S$ and $T$ parameters at 68% CL (dotted blue line) and 95% CL (full blue line). Here, light grey points are parameter points compatible with the constraints specified in Sec. 2, dark grey points additionally fulfil the SFOEWPT constraint $\xi_{p}>1$, and colored points fulfil the projected constraints on the EWPOs representatively obtained at a future FCC-$e$ for its $\sqrt{s}=m_{Z}$ phase. The $Z$-pole programme will efficiently single out a relatively small parameter space of the SFOEWPT-preferred parameter region. Panels 1(b) and 1(c) show the contours of the allowed mass spectra after successively applying the constraints. Panel 1(d) highlights the relation of this region in the context of the expected 125 GeV signal strength constraints obtainable at the HL-LHC at 68% CL (dotted red line) and 95% CL (dashed red line), displayed in the $\tan\beta$ versus $\cos(\beta-\alpha)$ plane.
The results for our parameter scan are presented in Fig. 1. As expected, parameter choices that are relatively far from the alignment limit can be efficiently constrained with a precision $Z$ pole programme at an $e^{+}e^{-}$ machine. Sensitivity here extends well beyond the sensitivity of a coupling analysis for the 125 GeV Higgs boson at the HL-LHC, which is highlighted in Fig. 1(d). [^5] To obtain the confidence intervals of the HL-LHC we performed a $\chi^{2}$ fit to test the projected values of the coupling modifiers to $b$- and $t$-quarks, $W$ and $Z$ bosons, gluons, and $\tau$ and $\mu$ leptons from Ref. CMS:2025hfp againts the 2HDM predictions from Eqs. (2) and (4). The requirement of a significant mass splitting between the heavy neutral scalar and pseudoscalar mass, $m_{H}$ and $m_{A}$, respectively, to achieve an SFOEWPT in the 2HDM (see also Dorsch:2014qja ; Dorsch:2016nrg ; Basler:2016obg ; Dorsch:2017nza ; Biekotter:2023eil ) stands in direct tension with an SM-like outcome of the oblique electroweak precision observables (EWPO) analysis for the expected FCC-ee precision, cf. Fig. 1(b). Points that still admit an SFOEWPT according to Eq. (1) are forced towards the alignment limit $\cos(\beta-\alpha)\simeq 0$, yet at relatively small mass scales, to modify the electroweak cross-over to an SFOEWPT, cf. Fig. 1(c). As already mentioned, signal strength analyses at the HL-LHC do not provide significant constraints for this parameter range. However, the mass scale of the exotic Higgs bosons falls into a range that is well accessible by ATLAS and CMS. We will turn to this in the next section.
@@ 3.2 Context with LHC Higgs Partner Searches
If the exotic states have masses above the $t\bar{t}$ threshold, the heavy scalars $H$ and $A$ predominantly decay into these final states. This holds for the largest part of the successful parameter points of our scan, detailed in Sec. 2. Both ATLAS and CMS are actively searching for heavy scalars in the $t\bar{t}$ channels CMS:2019pzc ; ATLAS:2024vxm ; CMS:2025dzq . It is well-established that interference effects, both between different Higgs resonances (signal–signal interference) Basler:2019nas ; Bahl:2025you and between the Higgs signal and the QCD background, can significantly distort the expected Breit–Wigner resonance shapes Gaemers:1984sj . In some cases, these effects can even result in a complete cancellation of the signal Jung:2015gta , or severely reduce experimental sensitivity, particularly when the net effect is a depletion rather than an excess of events.
@@
Figure 2: Interference-corrected (signal-signal as well as signal-background) cross sections compared to the pure signal process for LHC production of the uncharged heavy scalars, $H$ (blue points) and AA (red points), respectively, as a function of their respective mass, for the parameter points of the scan detailed in Sec. 2.
@@
We stress that the results presented here should not be understood as a ‘no-go theorem’ for the discovery of such states at the LHC. In fact, both ATLAS and CMS CMS:2025dzq ; ATLAS:2024vxm ; CMS:2019pzc incorporate interference effects in their likelihood analyses, yet with simplifying assumptions that make a direct comparison with our 2HDM scenario opaque. For instance, the recent ATLAS analysis ATLAS:2024vxm excludes parameter regions in the 2HDM type II with $m_{H}=m_{A}>300~\text{GeV}$ with $\tan\beta\lesssim 1.5-3$. It is worth highlighting that these exclusions are severely impacted by systematic uncertainties, as demonstrated by ATLAS.
The results for a 2HDM type I (as considered in this paper) would arguably be similar to those reported by ATLAS, because the phenomenology of all 2HDM types is very similar for $\tan\beta\sim 1$, and, moreover, the Higgs boson coupling to the top quark is universal for all types. Therefore, the ATLAS findings serve as a strong indicator that the LHC is gaining sensitivity in the mass region we consider. [^6] The observation of an excess near the $t\bar{t}$ threshold consistent with a “toponium” bound state by both ATLAS ATLAS:2025kvb and CMS CMS:2025kzt is evidence of this. However, precise non-relativistic QCD predictions of such a bound state will be necessary in the future for BSM searches in the $t\bar{t}$ channel in this mass range. Especially towards the HL-LHC phase, we can therefore expect these analyses to become increasingly sensitive, also to dip structures in the $m_{t\bar{t}}$ spectrum, which could then facilitate a discovery when the interference becomes large. To our knowledge, no HL-LHC extrapolation is currently available, but such analyses remain high-priority lines of BSM investigations with the potential to reveal new physics. In light of this, we will assume an overly pessimistic view that the LHC will not fully explore this region to show that an $e^{+}e^{-}$ machine can close this gap. Furthermore, other top-phillic search channels for neutral Higgs partners, such as top quark pair-associated and four top quark production, do not suffer from such sensitivity-altering effects. These, therefore, provide additional sensitivity; however, at a significantly reduced cross section and consequently limited sensitivity reach.
@@ 3.3.1 Elements of the Calculation
Figure 4: Production of the SM-like Higgs boson in association with a $Z$ boson at the FCC-$e$, in relation to the decoupling limit at $\sqrt{s}=240~\text{GeV}$ for the scan detailed in Sec. 2. Blue points denote leading-order predictions, which are all compatible with the expected FCC-ee sensitivity of hZhZ production of $\Delta\sigma/\sigma=0.31\%$ Selvaggi:2025kmd ; all points are compatible with an SM outcome of the FCC-ee $Z$-pole programme. Including radiative corrections (red points) modifies these predictions, creating the discovery potential at a lepton collider even when $Z$ pole observables are insensitive. The green region is the experimental sensitivity $\Delta\sigma/\sigma=0.31\%$. The decoupling cross section $\sigma_{\text{decoup}}$ is evaluated consistently at LO (blue points) and NLO (red points).
@@ 3.3.2 Higgs Production in $e^{+}e^{-}$ Collisions - Discrimination Beyond EWPOs
We are now ready to turn to the implications of our previous results for a precision investigation of associated Higgs production at an $e^{+}e^{-}$ collider. A Higgs machine will typically run at a centre-of-mass energy of $\sqrt{s}=240~\text{GeV}$ where the $e^{+}e^{-}\to hZ$ cross section has a maximum. The number of observables that are accessible in this instance is limited compared to hadron machines that probe a plethora of processes across a large range of energy scales. However, when viewed in the context of concrete model predictions, the overall sensitivity that can be expected at an FCC-ee of Selvaggi:2025kmd
$\frac{\Delta\sigma(e^{+}e^{-}\to hZ)}{\sigma(e^{+}e^{-}\to hZ)}=0.31\%\,$, (11)
@@
predominantly achieved from an analysis of the $Z$ boson recoil spectrum FCC:2018evy ; Azzurri:2021nmy ; TLEPDesignStudyWorkingGroup:2013myl , carries a very large potential for new physics discovery.
@@
The parameter region of our 2HDM scan that is consistent with the $Z$ pole programme is forced into the 2HDM alignment limit. Hence, it is not surprising that most of the parameter points are consistent with the uncertainty of Eq. (11) at the leading order. Consistency with S,TS,\,T forces the interactions of the SM-like Higgs boson into a region where the leading-order cross section is largely consistent with the SM (the 2HDM cross section in the decoupling limit is around $\sigma_{\mathrm{decoup}}=0.222$ pb at LO and $0.211$ pb at NLO), cf. Fig. 4.
@@
The requirement of an SFOEWPT, however, is incompatible with this decoupling limit (the latter approaches the SM for which the phase transition is a cross-over). Hence, we can expect quantum corrections to be relevant for the points that single out an SFOEWPT, also when they are aligned with the SM expectation of $Z$ pole measurements. Indeed, as shown in Fig. 4, the weak radiative corrections as detailed in Sec. 3.3.1, modify the 2HDM cross sections away from the decoupling limit (and hence, away from the SM). More importantly, the modifications are typically large enough for the $e^{+}e^{-}\to hZ$ programme to detect new physics based on the expectation of Eq. (11).
@@
It is worthwhile mentioning that the discrimination on the basis of the $hZ$ cross section is not in a one-to-one correspondence with the requirement of an SFOEWPT. It is the predictive power of renormalisability that drives experimentally detectable model correlations beyond leading order in a theoretically controlled way. An SFOEWPT is driven by the requirement of a relatively light exotic particle spectrum and a departure from SM-expected Higgs interactions. In any UV-complete scenario (and in particular for the 2HDM considered here), these requirements select a region of the parameter space that is, to a large extent, experimentally distinguishable from the SM in $hZ$ production at a 240 GeV lepton collider. The radiative corrections are not dominantly sensitive to the mass splittings themselves, but to their correlated effects in terms of inter-Higgs couplings and radiative corrections to mixing angles. These effects are two-loop suppressed for the $Z$ pole programme; $hZ$ production accessing these correlation modifications at one-loop level makes them large enough for detection within the expected experimental sensitivity. It is known that modifications of the Higgs self-coupling can be parametrically large Kanemura:2004mg ; Braathen:2019pxr ; Arco:2022lai ; Arco:2025pgx , including regions with SFOEWPTs Goncalves:2021egx ; Dorsch:2017nza ; Biekotter:2023eil ; Basler:2017uxn . These are formally two-loop suppressed effects, which could, however, be numerically relevant for the given sensitivity that can be achieved at a machine like FCC-ee. An analysis of such effects is only meaningful as part of a comprehensive two-loop investigation, which is beyond the scope of our present work.